# Designing and Evaluating Hint Generation Systems for Science Education

### v1

@@ 6.1.2. Designing for Learner Diversity and Goals
Aligning with learner goals. Every learner brings different objectives depending on the task and context (Hoque, 2016; Sönmez, 2017). Our findings reveal a similar learners’ preference variability, where some participants valued question-agnostic hints that guided them toward answers, while others who sought deeper understanding found these unhelpful. While providing context agnostic feedback to facilitate successful responses from learners improves short-term satisfaction, it hinders in development of critical thinking. While it is important to identify individual learning goals to provide the most suitable feedback, further caution is required to prevent over-reliance and over simplification. In Section 6.2, we discuss this tension of satisfaction with cognitive offloading in depth. Besides that, even identifying these learning goals isn’t a straightforward task, as prior work has shown, learners often lack the metacognitive awareness to articulate their own goals (Karpicke et al., 2009; Ku and Ho, 2010), and their learning preferences evolve over time (Gurpinar et al., 2011; Andrews and Tynan, 2011). Thus, tutoring systems must balance explicit input from learners with adaptive mechanisms that infer goals through interaction patterns. Metrics such as satisfaction ratings or longitudinal engagement may provide richer signals for tailoring support, pointing to the need for nuanced feedback frameworks that remain grounded in pedagogical values.
6.1.3. Beyond Cognition: Toward Holistic Learning Support
Need for mixed-initiative support. Prior work in intelligent tutoring shows that learners often delay or avoid seeking hints (Aleven and Koedinger, 2000), whereas human teachers frequently provide just-in-time guidance to prevent disengagement (Novak et al., 1999). Moreover, research has shown that low-performing students may lack the metacognitive skills necessary to make effective pedagogical decisions, such as when to request support (Aleven and Koedinger, 2000; Roll et al., 2014). Addressing this challenge, Ju et al. (2022) introduced a mixed-initiative tutoring framework using reinforcement learning, which successfully boosted the performance of low-achieving learners in a pretest–posttest study. While these results are promising, designing AI systems that can emulate the natural balance between learner autonomy and timely support remains a critical open challenge for future research.
@@ 6.1.3. Beyond Cognition: Toward Holistic Learning Support
Toward scaffolding self-regulated learning. While mixed-initiative frameworks can compensate for learners who struggle to recognize when support is needed, cultivating metacognitive skills themselves is essential for fostering critical thinking (Ku and Ho, 2010). Tutoring systems should therefore not only deliver assistance but also nurture learners’ capacity for self-regulation. Two pedagogical paradigms with strong empirical support offer promising directions: self-regulated learning (SRL) (Zimmerman, 1990, 2002; Prasad and Sane, 2024) and learning-by-teaching (LBT) (Frager and Stern, 1970). Both approaches have long-standing evidence of effectiveness in enhancing metacognition and deep learning (Allen and Feldman, 1973; Bargh and Schul, 1980; Chi et al., 2001; Cohen et al., 1982; Cheng, 2011; Bouchet et al., 2012), and recent advances in AI have opened new possibilities for embedding them in digital environments. For instance, Ge et al. (2025) demonstrate that LLM-assisted SRL can significantly improve regulation skills when scaffolded within a gamified environment. Likewise, Rogers et al. (2025) show how positioning LLMs as “tutees” enables learners to practice learning by teaching, thereby strengthening conceptual understanding as well as metacognitive reflection. Building on these insights, we envision future hint generation systems that extend beyond factual recall: systems that scaffold reflection, gradually fade support, and empower learners to assume increasing responsibility for their own learning. Such designs not only reduce risks of over-reliance but also promote the development of independent, critical thinkers.
6.2. Over-reliance and Cognitive Offloading
Recent years have seen rapid adoption of AI tools across everyday life including education, where students increasingly turn to AI assistance for help with assignments, writing, and problem-solving (Pitts et al., 2025a, b; Kanont et al., 2024; Lin, 2024). Prior work has documented this growing dependence, noting that learners often use these systems not only as aids but also as substitutes for their own cognitive effort (Adiyono et al., 2025). While such tools can provide immediate support, they also raise pressing concerns about their long-term effects on learning. Our findings echo this trend, as many participants preferring low-effort, context-agnostic hints that directly facilitated answering questions. This underscores a critical tension between short-term satisfaction and the risk of accumulating long-term cognitive debt. We contend that carefully designed hint generation systems could support learning without fostering over-reliance.
Several studies offer evidence that over-reliance on AI assistance may negatively affect cognitive engagement. Kosmyna et al. (2025) used electroencephalography (EEG) to measure neural activity during essay writing, comparing students who wrote independently against those who used LLM-based support. They found that AI-assisted writers exhibited markedly weaker neural connectivity patterns, suggesting reduced cognitive coupling during the task. Similarly, Gerlich (2025), through a mixed-method approach, found that younger learners reported higher dependence on AI tools and demonstrated lower critical thinking scores compared to older participants. Importantly, their findings also show that educational attainment correlates with stronger critical thinking skills regardless of AI use, pointing to both risks and potential mitigators of over-reliance. While these technologies can enhance writing proficiency and boost confidence, they also introduce risks of diminished creativity, over-reliance, and ethical concerns like plagiarism and systemic biases (Zhai et al., 2024).
Together, these findings highlight the urgency of designing educational technologies that support learners without undermining cognitive engagement. Rather than fully outsourcing intellectual effort, tutoring systems should aim to preserve opportunities for reflection, reasoning, and persistence. Recent research has taken steps toward this goal—for example, tools that emphasize explainability (Hoq et al., 2025), promote metacognitive reflection (Kumar et al., 2024; Yuan and Hu, 2024), or deliberately restrict direct answer-giving (Gupta et al., 2025). These approaches seek to balance the benefits of AI assistance with the preservation of deeper thinking.
Hint generation represents one such framework. By design, hints avoid revealing the complete answer, instead scaffolding learners’ problem-solving processes while encouraging persistence and self-efficacy. Learners often seek AI assistance when they are stuck, and unlike readily available AI systems that may provide full solutions, hint generation frameworks emphasize gradual support grounded in pedagogy principles, enabling learners to build knowledge rather than bypass it. Prior work in intelligent tutoring consistently shows that effective hints can increase motivation, improve self-efficacy, and foster a sense of accomplishment (Rogers et al., 2025; Mohammadi et al., 2025; VanLehn, 2011; Ma et al., 2014). Our findings reinforce this perspective: participants generally rated hints as helpful for answering questions, with several explicitly noting their value in promoting deeper thinking. For instance, P-22 described hints as ”hints encouraged to make associations and engage in reflection from a scientific perspective,” P-37 noted they ”encouraged active thinking,” and P-40 highlighted they ”encouraged me to look for other ways to say things.” Taken together with prior research, these results suggest that automatically generated hints can serve as a middle ground: providing support without fostering harmful dependence, and helping learners sustain their cognitive engagement in the presence of AI.
@@ 7. Limitations
Our study has several limitations that should be acknowledged when interpreting the findings. First, there are recruitment-related constraints. While our target demographic was high school students, ethical considerations around exposing minors to early-stage AI technology led us to recruit recent high school graduates instead. As a result, our participant pool consisted primarily of university students.
@@
Second, our study design reflects several methodological choices that constrain the scope of our findings. By focusing on factual questions, we primarily probed for remembering and understanding, rather than deeper levels of cognition such as application or analysis. This limits the generalizability of our insights to higher-order learning tasks. Moreover, our learner-driven setup assumes that participants possess sufficient metacognitive awareness to judge when to request hints. Prior research (Karpicke et al., 2009; Ku and Ho, 2010) suggests that this assumption does not always hold true, as many learners lack these skills, which in turn restricts the broader applicability of our proposed hint generation system. Additionally, our study also doesn’t probe the participants’ views of AI technologies, a factor that can affect the applicability of the results to broader learner population.
@@
Finally, there are limitations tied to the implementation of our hinting strategies. The dynamic hinting condition was unaware of the hint chain length, making it more difficult to plan hint sequencing and leading to a somewhat uneven comparison with the static condition. Furthermore, our system was agnostic to learners’ backgrounds and prior knowledge: while hints adapted superficially to student responses, they did not provide deeper personalization that could account for individual learning histories or subject expertise. Besides prior knowledge, our design also didn’t take into consideration affective dimensions like frustration, disengagement, or overconfidence, even though they are critical for learning (Mao and Li, 2010; Cunha-Pérez et al., 2018; Yadegaridehkordi et al., 2019; Hasan et al., 2020).
@@
Taken together, these limitations highlight important opportunities for future work. Recruiting participants closer to the intended learner demographic, extending evaluation to higher-order learning tasks, and developing adaptive systems that better account for learner diversity are all critical next steps. Addressing these gaps will help refine automatic hint generation into a more robust and equitable tool for supporting diverse educational contexts.
@@ 8. Conclusion
This paper examined the capabilities of large language models to generate and evaluate hints for science education. We benchmarked 18 open-source models using five automatic evaluation metrics grounded in learning sciences, each designed to capture a distinct attribute of effective feedback. We investigated two complementary hinting strategies: static hints, pre-generated for each problem, and dynamic hints, adapted to learners’ progress. Our quantitative study with 41 participants underscores both the promise and complexity of providing automated hints. Learners valued certain attributes that made hints effective, while also identifying aspects that hindered their learning experience. Building on these insights, we propose design principles for future systems that align with Bloom’s taxonomy of educational goals and emphasize learner-centered design. More broadly, we position automatic hint generation as a key research direction to address the growing challenge of cognitive debt among next-generation learners.
@@ A.1. Automatic Evaluation Metrics
(1) $\begin{split}InfoGain_{mean}(\mathbf{h})&=\frac{1}{k}\sum_{i=1}^{k}ROUGE_{recall}^{L}\bigl(LLM(q,h_{i}),a\bigr)\\ &\quad-ROUGE_{recall}^{L}\bigl(LLM_{QA}(q),a\bigr)\end{split}$
(2) $\begin{split}InfoGain_{comb}(\mathbf{h})&=ROUGE_{recall}^{L}\bigl(LLM(q,\mathbf{h}),a\bigr)\\ &\quad-ROUGE_{recall}^{L}\bigl(LLM_{QA}(q),a\bigr)\end{split}$
@@
Redundancy. Variation theory suggests that contrasting examples and perspectives help learners discern key features of a concept (Ling Lo, 2012). Lomibao and Ombay (2017)’s findings aligned with the theoretical argument of variation theory, the students exposed to repetition with variation to achieved significantly higher performance and improved knowledge retention in mathematical problem solving. To evaluate the diversity of perspectives in a chain-of-hints, we measure its self-referenced redundancy using the Sentence-BERT embeddings (Reimers and Gurevych, 2019) as proposed by Chen et al. (2021a) -
@@ A.2. Automatic Evaluation Results
Trends across model size. To formulate guidelines on what models to use for future research, we explore trends over three model families with multiple variants to observe how their hint generation capabilities varies across scale. We present the results for the largest family of models Qwen3 (Yang et al., 2025) in Figure LABEL:fig:scale_autoeval_main, and results for other two model families DeepSeek-R1 (Guo et al., 2025) and Gemma3 (Team et al., 2025) in Figure 14. We observe that the aggregate score increases with the model size, with static hints outperforming dynamic hints. This gap in performance decreases with the increase in number of parameters, indicating that the larger models are more capable at generating adaptive hints. The biggest contributor to this decrease in performance can be attributed to the answer leakage, where smaller models (<<4b parameters) have the exact answer string for over 20% of instances for Qwen3 and DeepSeek-R1 models. This leakage oversimplifies the problem, leading to an observable difference in the $InfoGain_{mean}$ metric, which measures the average information gain across a chain-of-hints. This effect is not as evident in the $InfoGain_{comb}$ metric, likely because it more holistically gauges the information gain of the entire chain, which even for larger models with less leakage, provide ample information to help the information gain evaluator LM. Smaller models are also not very expressive, with more redundant hint sequences compared to larger models, that provide more diverse perspectives due to their better memorization and reasoning capabilities. While there’s a definitive trend of improved hint quality with an increase in size, for a simpler task of factual scientific question answering, the performance plateaus after a certain size, with models between 12-14b parameters performing sufficiently well off-the-shelf. For a resource constrained domain of education, we believe further finetuning these medium size models to be the most pragmatic next direction.
@@ Appendix B Prompts
We provide the prompts we used throughout our study in this section. Figures 10 and 11 describe the static and dynamic hint generation prompts we used for both automatic evaluation benchmarking and the human evaluation study. Figure 12 contains the prompt we used to measure the LLM-based answer leakage in $leakage_{LLM}$ evaluation metric (using Gemma3-27B model (Team et al., 2025)). Figure 13 comprises of the answer assessment prompt we used to resolve fuzzy matching of participants’ attempted answer with the ground truth answer (using OpenAI’s o4-mini model). The answer assessment was preceded by an exact string match to increase the assessment performance due to the non-deterministic nature of LLM responses. While the figures present the prompt in two parts, a system prompt and the user prompt, in practice, we treated them as a single string to ensure a fair comparison across all models. We leave the finetuning of these prompts to future works.

### v2

@@ 6.2.2. Designing for Learner Diversity and Goals
Aligning with learner goals. Every learner brings different objectives depending on the task and context (Hoque, 2016; Sönmez, 2017). Our findings reveal a similar learners’ preference variability, where some participants valued question-agnostic hints that guided them toward answers, while others who sought deeper understanding found these unhelpful. While providing context agnostic feedback to facilitate successful responses from learners improves short-term satisfaction, it hinders in development of critical thinking. While it is important to identify individual learning goals to provide the most suitable feedback, further caution is required to prevent over-reliance and over simplification. Besides that, even identifying these learning goals isn’t a straightforward task, as prior work has shown, learners often lack the metacognitive awareness to articulate their own goals (Karpicke et al., 2009; Ku and Ho, 2010), and their learning preferences evolve over time (Gurpinar et al., 2011; Andrews and Tynan, 2011). Thus, tutoring systems must balance explicit input from learners with adaptive mechanisms that infer goals through interaction patterns. Metrics such as satisfaction ratings or longitudinal engagement may provide richer signals for tailoring support, pointing to the need for nuanced feedback frameworks that remain grounded in pedagogical values.
@@
Toward scaffolding self-regulated learning. We found that even the hint generation systems explicitly designed to promote cognitive engagement can still exhibit a tendency to produce overly answer-oriented hints (Section 6.1). Prolonged exposure to such systems may inadvertently reinforce overreliance, amplifying the cognitive offloading they were intended to mitigate. Therefore, cultivating metacognitive skills is essential for fostering critical thinking (Ku and Ho, 2010). Tutoring systems should not only deliver assistance but also nurture learners’ capacity for self-regulation. Two pedagogical paradigms with strong empirical support offer promising directions: self-regulated learning (SRL) (Zimmerman, 1990, 2002; Prasad and Sane, 2024) and learning-by-teaching (LBT) (Frager and Stern, 1970). Both approaches have long-standing evidence of effectiveness in enhancing metacognition and deep learning (Allen and Feldman, 1973; Bargh and Schul, 1980; Chi et al., 2001; Cohen et al., 1982; Cheng, 2011; Bouchet et al., 2012), and recent advances in AI have opened new possibilities for embedding them in digital environments. For instance, Ge et al. (2025) demonstrate that LLM-assisted SRL can significantly improve regulation skills when scaffolded within a gamified environment. Likewise, Rogers et al. (2025) show how positioning LLMs as “tutees” enables learners to practice learning by teaching, thereby strengthening conceptual understanding as well as metacognitive reflection. Building on these insights, we envision future hint generation systems that extend beyond factual recall: systems that scaffold reflection, gradually fade support, and empower learners to assume increasing responsibility for their own learning. Such designs not only reduce risks of over-reliance but also promote the development of independent, critical thinkers.
@@ 7. Limitations
Our study has several limitations that should be acknowledged when interpreting the findings. First, there are recruitment-related constraints. While our target demographic was high school students, ethical considerations around exposing minors to early-stage AI technology led us to recruit recent high school graduates instead. As a result, our participant pool consisted primarily of university students. Rather than treating the study as a direct evaluation of high school science learning, we position it as an exploratory investigation of hint–learner interaction dynamics in a controlled scientific question-answering setting.
@@
Second, our study design reflects several methodological choices that constrain the scope of our findings. By focusing on factual questions, we primarily probed for remembering and understanding, rather than deeper levels of cognition such as application or analysis. This limits the generalizability of our insights to higher-order learning tasks. Additionally, our study also doesn’t probe the participants’ views of AI technologies, a factor that can affect the applicability of the results to broader learner population.
@@
Third, there are limitations tied to the implementation of our hinting strategies. Our work does not conduct an exhaustive search in the prompt space for hint generation, and uses same prompts across all models. Due to lack of prior interaction data, we limit the investigation to zero-shot prompts, that lack in-context supervision to improve performance (Brown et al., 2020). For a fair assessment and comparison across several models, the hint generation prompts are quite simple. We leave the investigation of using behavioral signals, learning feedback loops and expert-written instructions in hint generation prompts to future work. Additionally, the dynamic hinting condition was unaware of the hint chain length, making it more difficult to plan hint sequencing and leading to a somewhat uneven comparison with the static condition. Furthermore, our system was agnostic to learners’ backgrounds and prior knowledge: while hints adapted superficially to student responses, they did not provide deeper personalization that could account for individual learning histories or subject expertise. Besides prior knowledge, our design also didn’t explicitly take into consideration affective dimensions like frustration, disengagement, or overconfidence, even though they are critical for learning (Mao and Li, 2010; Cunha-Pérez et al., 2018; Yadegaridehkordi et al., 2019; Hasan et al., 2020).
Fourth, our study relies on the SciQ dataset (Welbl et al., 2017a) as the source of question–answer pairs. Although the dataset is derived from high school science materials, it was originally constructed for benchmarking question-answering systems rather than for evaluating pedagogical scaffolding. As a result, the factual and decontextualized nature of SciQ limits the ecological validity of our setting and may not fully capture the complexity of real instructional interactions. Moreover, our automatic evaluation pipeline for answer assessment and detecting leakage relies on LLM-based judgments rather than human-annotated ground truth. This introduces potential self-consistency bias, reproducibility concerns, and misalignment with human perceptions of hint quality. While we compliment this by conducting a user study, the dependence on LLM-based automatic metrics to select the model for user study remains a limitation. Future work should develop pedagogy-grounded benchmarks and human-validated evaluation protocols tailored specifically for assessing hinting and scaffolding behaviors.
@@
Taken together, these limitations highlight important opportunities for future work. Recruiting learners closer to the intended demographic, expanding beyond factoid questions to incorporate higher-order reasoning tasks, and designing interfaces that support diverse help-seeking behaviors will improve ecological validity. Just as critically, moving beyond QA-oriented datasets like SciQ and reducing reliance on LLM-based self-evaluation will be necessary to build more reliable, pedagogy-grounded benchmarks for assessing hint quality. Addressing these gaps will help refine automatic hint generation into a more robust and equitable tool for supporting diverse educational contexts.
@@ 8. Conclusion
This paper examined the capabilities of large language models to generate and evaluate hints for science education. We benchmarked 18 open-source models using five automatic evaluation metrics grounded in learning sciences, each designed to capture a distinct attribute of effective feedback. We investigated two complementary hinting strategies: static hints, pre-generated for each problem, and dynamic hints, adapted to learners’ progress. Our quantitative study with 41 participants underscores both the promise and complexity of providing automated hints. Participants valued certain attributes that made hints effective, while also identifying aspects that hindered their learning experience. Building on these insights, we propose design principles for future systems that align with Bloom’s taxonomy of educational goals and emphasize learner-centered design. More broadly, we position automatic hint generation as a key research direction to address the growing challenge of cognitive debt among next-generation learners.
@@ A.1. Automatic Evaluation Metrics
(1) $\begin{split}InfoGain_{mean}(\mathbf{h})&=\frac{1}{k}\sum_{i=1}^{k}ROUGE_{recall}^{L}\bigl(LLM_{QA}(q,h_{i}),a\bigr)\\ &\quad-ROUGE_{recall}^{L}\bigl(LLM_{QA}(q),a\bigr)\end{split}$
(2) $\begin{split}InfoGain_{comb}(\mathbf{h})&=ROUGE_{recall}^{L}\bigl(LLM_{QA}(q,\mathbf{h}),a\bigr)\\ &\quad-ROUGE_{recall}^{L}\bigl(LLM_{QA}(q),a\bigr)\end{split}$
@@
Redundancy. Variation theory suggests that contrasting examples and perspectives help learners discern key features of a concept (Ling Lo, 2012). Lomibao and Ombay (2017)’s findings aligned with the theoretical argument of variation theory, the students exposed to repetition with variation to achieved significantly higher performance and improved knowledge retention in mathematical problem solving. To evaluate the diversity of perspectives in a chain-of-hints, we measure its self-referenced redundancy using the Sentence-BERT embeddings (Reimers and Gurevych, 2019) as proposed by Chen et al. (2021b) -
@@ A.2. Automatic Evaluation Results
Trends across model size. To formulate guidelines on what models to use for future research, we explore trends over three model families with multiple variants to observe how their hint generation capabilities varies across scale. We present the results for the three model families Qwen3 (Yang et al., 2025), DeepSeek-R1 (Guo et al., 2025) and Gemma3 (Team et al., 2025) in Figure 15. We observe that the aggregate score increases with the model size, with static hints outperforming dynamic hints. This gap in performance decreases with the increase in number of parameters, indicating that the larger models are more capable at generating adaptive hints. The biggest contributor to this decrease in performance can be attributed to the answer leakage, where smaller models (<<4b parameters) have the exact answer string for over 20% of instances for Qwen3 and DeepSeek-R1 models. This leakage oversimplifies the problem, leading to an observable difference in the $InfoGain_{mean}$ metric, which measures the average information gain across a chain-of-hints. This effect is not as evident in the $InfoGain_{comb}$ metric, likely because it more holistically gauges the information gain of the entire chain, which even for larger models with less leakage, provide ample information to help the information gain evaluator LM. Smaller models are also not very expressive, with more redundant hint sequences compared to larger models, that provide more diverse perspectives due to their better memorization and reasoning capabilities. While there’s a definitive trend of improved hint quality with an increase in size, for a simpler task of factual scientific question answering, the performance plateaus after a certain size, with models between 12-14b parameters performing sufficiently well off-the-shelf. For a resource constrained domain of education, we believe further finetuning these medium size models to be the most pragmatic next direction.
@@ Appendix B Prompts
We provide the prompts we used throughout our study in this section. Figures 11 and 12 describe the static and dynamic hint generation prompts we used for both automatic evaluation benchmarking and the human evaluation study. Figure 13 contains the prompt we used to measure the LLM-based answer leakage in $leakage_{LLM}$ evaluation metric (using Gemma3-27B model (Team et al., 2025)). Figure 14 comprises of the answer assessment prompt we used to resolve fuzzy matching of participants’ attempted answer with the ground truth answer (using OpenAI’s o4-mini model). The answer assessment was preceded by an exact string match to increase the assessment performance due to the non-deterministic nature of LLM responses. While the figures present the prompt in two parts, a system prompt and the user prompt, in practice, we treated them as a single string to ensure a fair comparison across all models. We leave the finetuning of these prompts to future works.